# A positive density of elliptic curves are diophantine stable in certain Galois Extensions

### v1

@@ Abstract.
Let $p\in\{3,5\}$ and consider a cyclic $p$-extension $L/\mathbb{Q}$. We show that there exists an effective positive density of elliptic curves $E$ defined over $\mathbb{Q}$, ordered by height, that are diophantine stable in $L$.
@@ 1.1. Background and motivation
Let $E$ be an elliptic curve over a number field $K$ and $L$ be a finite extension of $K$. Following Mazur and Rubin [MR18], $E$ is said to be diophantine stable in $L$ if the natural inclusion induces an equality $E(K)=E(L)$. Mazur and Rubin show that if all $\bar{K}$-automorphisms of $E$ are defined over $K$, then there exist infinitely many prime-power cyclic extensions $L/K$ in which $E$ is diophantine stable. We refer to [MR18, Theorem 1.2 and Theorem 1.3] for further details. For elliptic curves over $\mathbb{Q}$, there are precise heuristics for the growth of the rank, due to David, Fearnley and Kisilevsky [DFK07]. These heuristics are based on the relationship between modular symbols and the special values of the Hasse-Weil $L$-function $\mathscr{L}_{E}(s,L)$. More precisely, let $L/\mathbb{Q}$ be a Galois extension with cyclic Galois group $G=\operatorname{Gal}(L/\mathbb{Q})$. Then the Hasse-Weil $L$-function of $E$ over $L$ decomposes as follows $\mathscr{L}_{E}(s,L)=\prod_{\chi}\mathscr{L}_{E}(s,\chi)$, where $\chi$ ranges over all Dirichlet characters $\chi:\widehat{G}\rightarrow\mathbb{C}^{\times}$. The Birch and Swinnerton–Dyer conjecture predicts that
@@
Given a character $\chi$, set $\operatorname{cond}(\chi)$ to denote its conductor. One has the following conjecture which predicts that the rank rarely jumps in prime cyclic extensions.
Conjecture 1.1 (Conjecture 1.2 of [DFK07]).
Let $p$ be an odd prime and fix an elliptic curve $E_{/\mathbb{Q}}$. Setting
$N_{E,p}(X):=\#\{\chi\text{ of order }p\mid\operatorname{cond}(\chi)\leq X\text{ and }\mathscr{L}_{E}(1,\chi)=0\},$
one has the following predictions.
  1. (1)
    If $p=3$, then as $X\to\infty$,
    $\log N_{E,p}(X)\sim\frac{1}{2}\log X.$
  2. (2)
    If $p=5$, then as $X\to\infty$, $N_{E,p}(X)$ is unbounded but $N_{E,p}(X)\ll X^{\epsilon}$ for any $\epsilon>0$.
  3. (3)
    If $p\geq 7$, then $N_{E,p}(X)$ is bounded.
In particular, the conjecture predicts that when $p\geq 7$, there are only finitely many $p$-cyclic extensions in which the rank increases. There has been an increasing interest in the study of diophantine stability and rank growth of a given elliptic curve in families of number field extensions, see for instance [MR08, LOT21, SW23, BKR24, BRY24]. Such considerations also motivate conjectures about the distribution of modular symbols associated with a given elliptic curves, cf. [MR23].
@@ 1.2. Main result
    for all $\ell\in Z$, there exists an elliptic curve $\mathbb{E}$ over $\mathbb{F}_{\ell}$, such that $\mathbb{E}(\mathbb{F}_{\ell})[p]=0$ and $\mathbb{E}^{-1}(\mathbb{F}_{\ell})[p]=0$. Here, $\mathbb{E}^{-1}$ is the quadratic twist of $\mathbb{E}$ by $-1$.
@@
  • The results rely on the techniques of Bhargava and Shankar [BS15a, BS15b, BS13], who study the average size of the $p$-Selmer group of elliptic curves $E_{/\mathbb{Q}}$. This is only done for primes $p=2,3,5$, however, the case when $p=2$ is omitted since our methods do not apply to this case.
  • The condition (2) can be dealt with via explicit calculation (cf. the code in Remark 4.7). Indeed, when $p=5$ and $\ell$ is large enough, the condition (2) is shown to be be satisfied using results of Howe [How93]. We refer to Lemma 4.6 for further details.
  • An explicit lower density is obtained for the curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$. This expression is quite fascinating and involves locally defined invariants associated to the extension $L/\mathbb{Q}$, see the statement of Theorem 3.8 for details. For instance, when $p=3$ and $L=\mathbb{Q}(\mu_{7})^{+}$, the lower density of elliptic curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$ is
    $\geq\frac{1}{4}\times\frac{1}{2^{21}}\times\left(1-\frac{1}{3}\right)\times\frac{12}{49}\times\prod_{\ell\neq 2,3,7}\left(1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}\right).$
@@ 1.3. Organization
Including the introduction, the article consists of four sections. In section 2, we study the stability of Selmer groups in prime cyclic extensions. In greater detail, let $p$ be an odd prime number and $L/K$ be an extension with $\operatorname{Gal}(L/K)\simeq\mathbb{Z}/p\mathbb{Z}$. Suppose that the $p$-Selmer group of $E$ over $K$ is 00. Then if certain additional local condition are satisfied for $E_{/K}$, then, it follows that the $p$-Selmer group of $E$ over $L$ is also 00. The Galois cohomological arguments presented in this section are essentially due to Česnavičius [Č17] and Brau [Bra14]. Section 3 is preparatory in nature, and delineates the strategy that we take to prove our main result. We introduce both local and global density conditions on our families of elliptic curves that will play a role in our sieve theoretic arguments. Section 4 is devoted to the proof of our main result. In this section, we describe a suitable large family of elliptic curves defined by local conditions and show that a positive proportion of elliptic curves in this family are indeed diophantine stable in $L$. The analysis leads to an explicit lower bound for the density of elliptic curves that are diophantine stable in our fixed extension $L$. This is precisely the step in which the aforementioned results of Bhargava and Shankar are applied.
@@ 2. Selmer groups and diophantine stability
In this section, we review the necessary background and notation required for our examination of elliptic curves over number fields. Some of the results in this section essentially follow arguments due to Česnavičius [Č17] and Brau [Bra14]. Throughout, we denote $K$ as a number field. We use $\Omega(K)$ (resp. $\Omega_{\infty}(K)$) to represent the set of finite primes (resp. archimedean places) of $K$. Given a prime number $\ell$, $\Omega_{\ell}(K)$ denotes the primes in $\Omega(K)$ that lie above $\ell$. Let $\overline{K}$ be an algebraic closure of $K$, and $\operatorname{G}_{K}$ denote the absolute Galois group $\operatorname{Gal}(\overline{K}/K)$. For each prime $v\in\Omega(K)$, we select an embedding $\iota_{v}:\overline{K}\hookrightarrow\overline{K}_{v}$. Setting $\operatorname{G}_{K_{v}}:=\operatorname{Gal}(\overline{K}_{v}/K_{v})$, the inclusion $\iota_{v}$ induces a Galois group inclusion $\iota_{v}^{*}:\operatorname{G}_{K}\hookrightarrow\operatorname{G}_{K_{v}}$.
@@ 2.1. Selmer groups associated to elliptic curves
$\operatorname{Sel}_{p}(E/K)=\ker\left(H^{1}(K,E[p^{n}])\to\bigoplus_{v}H^{1}(K_{v},E)[p^{n}]\right),$
@@
$0\rightarrow E(K)/p^{n}E(K)\rightarrow\operatorname{Sel}_{p^{n}}(E/K)\rightarrow\Sh(E/K)[p^{n}]\rightarrow 0$
@@ 2.2. Growth and stability of Selmer groups
Let $E_{/\mathbb{Q}}$ be an elliptic curve and $K$ be a number field. Let $p$ be an odd prime number. Fix a finite Galois extension $L/K$ for which $G:=\operatorname{Gal}(L/K)$ is a $p$-group and assume that $G\simeq\mathbb{Z}/p\mathbb{Z}$. Set $k_{v}$ to denote the residue field of $K$ at $v$. If $v$ is a prime of good reduction, we set $\widetilde{E}(k_{v})$ to be the group of $k_{v}$-rational points of the mod-$v$ reduction of $E$. Let $c_{v}(E/K)$ be the Tamagawa number of $E$ at $v$. Throughout, we impose the following assumption.
@@
    The Selmer group $\operatorname{Sel}_{p}(E/K)=0$. In particular, this implies that the rank of $E(K)$ is 00 and both $E(K)$ and $\Sh(E/K)$ have no nontrivial $p$-torsion.
@@
$\alpha:\operatorname{Sel}_{p}(E/K)\rightarrow\operatorname{Sel}_{p}(E/L)^{G}$
@@
$\gamma_{v}:H^{1}(K_{v},E)[p]\rightarrow\left(\bigoplus_{w|v}H^{1}(L_{w},E)[p]\right)^{G},$
@@
The result follows via an easy application of Nakayama’s lemma. For further details, cf. [MR24, Lemma 4.2]. ∎
@@
(2.2) $0=\operatorname{ker}(\alpha)\rightarrow\operatorname{ker}(\beta)\rightarrow\operatorname{ker}(\gamma)\rightarrow\operatorname{cok}(\alpha)\rightarrow\operatorname{cok}(\beta)\rightarrow\operatorname{cok}(\gamma).$
@@
[…]
it follows that $\beta$ is an isomorphism. This implies that $\operatorname{ker}\gamma$ is isomorphic to the $\operatorname{cok}\alpha$. Since the maps $\gamma_{v}$ are injective, so is $\gamma$. Hence, $\operatorname{Sel}_{p}(E/L)^{G}=0$. The result then follows from Lemma 2.2, which asserts that $\operatorname{Sel}_{p}(E/L)=0$. ∎
@@
  • $v$ is unramified in $L$ and $E$ has nonsplit multiplicative or additive reduction at $v$,
@@
It follows from the definition of $S$ that $v$ must be totally ramified in $L$. Recall that Assumption 2.1 requires that $\widetilde{E}(k_{v})[p]=0$. The result then follows from [PR24, Proposition 2.8]. ∎
@@
First we consider the case in which $E$ has nonsplit multiplicative reduction or additive reduction at $v$. In this case, the result follows from [Bra14, Proposition 5.8]. On the other hand, consider the case in which $E$ has split multiplicative reduction. Note that in light of Assumption 2.1, $v$ must be unramified in $L$ and $p\nmid c_{v}(E/K)$. Then, the result follows from [Bra14, Proposition 5.7, (ii)]. ∎
@@
  • First, assume that $v$ is unramified in $L$ and that $E$ has good ordinary reduction at $v$. In this case, $\gamma_{v}$ is injective by [Bra14, Proposition 5.9, (ii)].
  • Next consider the case when $v$ is ramified in $L$ and $E$ has good ordinary reduction at $v$. In this case, it has been assumed that $\widetilde{E}(k_{v})[p]=0$, by Assumption 2.1. In this case, $\gamma_{v}$ is injective, according to [Bra14, Proposition 5.9, (i)].
  • Consider the case when $v$ is unramified in $L$ and $E$ has good supersingular reduction at $v$. Then, $\gamma_{v}$ is injective by [Bra14, Proposition 5.10, (ii)].
@@ 3. Asymptotics for elliptic curves
$h_{A,B}=h\left(E_{A,B}\right):=\operatorname{max}\{|A|^{3},|B|^{2}\}.$
@@
$\mathcal{C}(X):=\{(A,B)\in\mathcal{C}:h\left(E_{A,B}\right)\leq X\}.$
@@
Let $\mathcal{C}(X)$ be given as above. One has the following asymptotic estimate for the growth of #$\#\mathcal{C}(X)$, as $X$ goes to $\infty$:
#$\#\mathcal{C}(X)=\frac{4X^{5/6}}{\zeta(10)}+O\left(\sqrt{X}\right).$
@@
Any set of isomorphism classes of elliptic curves over $\mathbb{Q}$ can be identified with a subset $S$ of $\mathcal{C}$. Given $X>0$, set $S(X):=S\mathcal{A}p\mathcal{C}(X)$. The density of $S$ (if it exists) is defined as follows
(3.1) $\mathfrak{d}(S):=\lim_{X\rightarrow\infty}\frac{\#S(X)}{\#\mathcal{C}(X)}.$
[…]
@@
The density of elliptic curves $E_{/\mathbb{Q}}$ with rank 00 (resp. $1$) is $1/2$ (resp. $1/2$).
The conjecture thus also predicts that the density of elliptic curves with rank $\geq 2$ is 00. Although this conjecture is open, significant progress was made in seminal work of Bhargava and Shankar [BS15a, BS15b, BS13]. The idea is to relate Selmer groups of elliptic curves to integral forms. The integral forms are lattice points in certain fundamental regions in Euclidean space and techniques from the geometry of numbers are applied to study their distribution. Given a prime $p$, it is expected that the average size of the $p$-Selmer group of $E_{/\mathbb{Q}}$ is $(p+1)$. This has been proven to be the case for primes $p\leq 5$. We refer to work of Poonen and Rains [PR12] for predictions on the average size of Selmer groups and Tate–Shafarevich groups.
In this section, we fix a Galois extension $L/\mathbb{Q}$ with Galois group $\operatorname{Gal}(L/\mathbb{Q})\simeq\mathbb{Z}/p\mathbb{Z}$, where $p\in\{3,5\}$. Let $E_{/\mathbb{Q}}$ be an elliptic curve. Following Mazur and Rubin [MR18], $E$ is said to be diophantine stable in $L$ if the natural inclusion yields an equality $E(\mathbb{Q})=E(L)$. We assume throughout that $2$ and $p$ are unramified in $L$, and show that there is an effective positive density of elliptic curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$. This is done by showing that

### v2

# A positive density of elliptic curves is diophantine stable in certain Galois Extensions

@@ Abstract.
Let $p\in\{3,5\}$ and consider a cyclic $p$-extension $L/\mathbb{Q}$. We show that there exists an effective positive density of elliptic curves $E$ defined over $\mathbb{Q}$, ordered by height, that is diophantine stable in $L$.
Key words and phrases:
Diophantine stability, arithmetic statistics, elliptic curves, rank stability
@@ 1.1. Background and motivation
Let $E$ be an elliptic curve over $\mathbb{Q}$ and $L$ be a number field. Following Mazur and Rubin [Mazurrubindiophantine], $E$ is said to be diophantine stable if the natural inclusion induces an equality $E(L)=E(\mathbb{Q})$. Mazur and Rubin show that if all $\bar{\mathbb{Q}}$-automorphisms of $E$ are defined over $\mathbb{Q}$, then there exist infinitely many prime-power cyclic extensions $L/\mathbb{Q}$ in which $E$ is diophantine stable. We refer to [Mazurrubindiophantine, Theorem 1.2 and Theorem 1.3] for further details.
Assume that $L/\mathbb{Q}$ is Galois with cyclic Galois group $G=\operatorname{Gal}(L/\mathbb{Q})$. Then the Hasse-Weil $L$-function of $E$ over $L$ decomposes as follows $\mathscr{L}_{E}(s,L)=\prod_{\chi}\mathscr{L}_{E}(s,\chi)$, where $\chi$ ranges over all Dirichlet characters $\chi\mathrel{\mathop{\ordinarycolon}}\widehat{G}\rightarrow\mathbb{C}^{\times}$. The Birch and Swinnerton–Dyer conjecture predicts that
@@
There are heuristics for the growth and stability of ranks of elliptic curves $E$ in cyclic extensions, which are based on the relationship between modular symbols and the values of $\operatorname{ord}_{s=1}\mathscr{L}_{E}(s,\chi)$ above. For further details, we refer to [Mazurrubinmodsymbols].
There has been an increasing interest in the study of diophantine stability and rank growth of a given elliptic curve in families of number field extensions, see for instance [MazurRubingrowthofSelmer, Mazurrubingrowthnonabelian, Oliverthorne, ShnidmanWeiss, BKR, Berg]. Moreover, the study of such rank stability questions has close connections to Hilbert’s tenth problem, see [MazurRubinH10Inventiones]. These connections have recently been studied with rejuvenated interest by Garcia–Fritz and Pasten [garcia-fritzpasten] and Kundu–Lei–Sprung [KLS].
@@ 1.2. Main result
    for all $\ell\in Z$, there exists an elliptic curve $\mathbb{E}$ over $\mathbb{F}_{\ell}$, such that $\mathbb{E}(\mathbb{F}_{\ell})[p]=0$ and $\mathbb{E}^{-1}(\mathbb{F}_{\ell})[p]=0$. Here, $\mathbb{E}^{-1}$ is the quadratic twist of $\mathbb{E}$ by 1-1.
@@
  • The results rely on the techniques of Bhargava and Shankar [BhargavaShankarbinary, bhargavashankar15, bhargavashankar5], who study the average size of the $p$-Selmer group of elliptic curves $E_{/\mathbb{Q}}$. This is only done for primes $p=2,3,5$, however, the case when $p=2$ is omitted since our methods do not apply to this case.
  • The condition (2) can be dealt with via explicit calculation (cf. the code in Remark 4.7). Indeed, when $p=5$ and $\ell$ is large enough, the condition (2) is shown to be be satisfied using results of Howe [Howe]. We refer to Lemma 4.6 for further details.
  • An explicit lower density is obtained for the curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$. This expression is quite fascinating and involves locally defined invariants associated to the extension $L/\mathbb{Q}$, see the statement of Theorem 3.8 for details. For instance, when $p=3$ and $L\subset\mathbb{Q}(\mu_{31})$ is the cubic subfield, the lower density of elliptic curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$ is
    $\geq\frac{1}{4}\times\frac{1}{2^{21}}\times\left(1-\frac{1}{3}\right)\times\frac{585}{961}\times\prod_{\ell\neq 2,3,31}\left(1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}\right).$
@@ 1.3. Organization
Including the introduction, the article consists of four sections. In section 2, we study the stability of Selmer groups in prime cyclic extensions. In greater detail, let $p$ be an odd prime number and $L/K$ be an extension with $\operatorname{Gal}(L/K)\simeq\mathbb{Z}/p\mathbb{Z}$. Suppose that the $p$-Selmer group of $E$ over $K$ is 0. Then if certain additional local condition are satisfied for $E_{/K}$, then, it follows that the $p$-Selmer group of $E$ over $L$ is also 0. The Galois cohomological arguments presented in this section are essentially due to Česnavičius [navicius] and Brau [brau2014selmer]. Section 3 is preparatory in nature, and delineates the strategy that we take to prove our main result. We introduce both local and global density conditions on our families of elliptic curves that will play a role in our sieve theoretic arguments. Section 4 is devoted to the proof of our main result. In this section, we describe a suitable large family of elliptic curves defined by local conditions and show that a positive proportion of elliptic curves in this family are indeed diophantine stable in $L$. The analysis leads to an explicit lower bound for the density of elliptic curves that are diophantine stable in our fixed extension $L$. This is precisely the step in which the aforementioned results of Bhargava and Shankar are applied.
@@
Acknowledgment
We thank the anonymous referee for the helpful report.
@@ 2. Selmer groups and diophantine stability
In this section, we review the necessary background and notation required for our examination of elliptic curves over number fields. Some of the results in this section extend arguments due to Česnavičius [navicius] and the unpublished work of Brau [brau2014selmer]. Throughout, we denote by $K$ a number field and use $\Omega(K)$ (resp. $\Omega_{\infty}(K)$) to represent the set of finite primes (resp. archimedean places) of $K$. Given a prime number $\ell$, $\Omega_{\ell}(K)$ denotes the primes in $\Omega(K)$ that lie above $\ell$. Let $\overline{K}$ be an algebraic closure of $K$, and $\operatorname{G}_{K}$ denote the absolute Galois group $\operatorname{Gal}(\overline{K}/K)$. For each prime $v\in\Omega(K)$, we select an embedding $\iota_{v}\mathrel{\mathop{\ordinarycolon}}\overline{K}\hookrightarrow\overline{K}_{v}$. Setting $\operatorname{G}_{K_{v}}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\overline{K}_{v}/K_{v})$, the inclusion $\iota_{v}$ induces a Galois group inclusion $\iota_{v}^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{K}\hookrightarrow\operatorname{G}_{K_{v}}$.
@@ 2.1. Selmer groups associated to elliptic curves
$\operatorname{Sel}_{p^{n}}(E/K)=\ker\left(H^{1}(K,E[p^{n}])\to\bigoplus_{v}H^{1}(K_{v},E)[p^{n}]\right),$
@@
$0\rightarrow E(K)/p^{n}E(K)\rightarrow\operatorname{Sel}_{p^{n}}(E/K)\rightarrow\Sh(E/K)[p^{n}]\rightarrow 0.$
@@ 2.2. Growth and stability of Selmer groups
Let $E_{/\mathbb{Q}}$ be an elliptic curve and $K$ be a number field. Let $p$ be an odd prime number. Fix a finite Galois extension $L/K$ for which $G\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(L/K)$ is a $p$-group and assume that $G\simeq\mathbb{Z}/p\mathbb{Z}$. Set $k_{v}$ to denote the residue field of $K$ at $v$. If $v$ is a prime of good reduction, we set $\widetilde{E}(k_{v})$ to be the group of $k_{v}$-rational points of the mod-$v$ reduction of $E$. Let $c_{v}(E/K)$ be the Tamagawa number of $E$ at $v$.
@@
    The Selmer group $\operatorname{Sel}_{p}(E/K)=0$. In particular, this implies that the rank of $E(K)$ is 0 and both $E(K)$ and $\Sh(E/K)$ have no nontrivial $p$-torsion.
@@
$\alpha\mathrel{\mathop{\ordinarycolon}}\operatorname{Sel}_{p}(E/K)\rightarrow\operatorname{Sel}_{p}(E/L)^{G}$
@@
$\gamma_{v}\mathrel{\mathop{\ordinarycolon}}H^{1}(K_{v},E)[p]\rightarrow\left(\bigoplus_{w|v}H^{1}(L_{w},E)[p]\right)^{G},$
@@
The result follows via an easy application of Nakayama’s lemma. For further details, cf. [muller2024hilbert, Lemma 4.2]. ∎
@@
(2.2) $0=\operatorname{ker}(\alpha)\rightarrow\operatorname{ker}(\beta)\rightarrow\operatorname{ker}(\gamma)\rightarrow\operatorname{coker}(\alpha)\rightarrow\operatorname{coker}(\beta)\rightarrow\operatorname{coker}(\gamma).$
@@
$0\rightarrow H^{1}(L/K,E(L)[p])\rightarrow H^{1}(K_{S}/K,E[p])\xrightarrow{\beta}H^{1}(K_{S}/L,E[p])^{G}\rightarrow H^{2}(L/K,E(L)[p])$
it follows that $\beta$ is an isomorphism. This implies that $\operatorname{ker}\gamma$ is isomorphic to the $\operatorname{coker}\alpha$. Since the maps $\gamma_{v}$ are injective, so is $\gamma$. Hence, $\operatorname{Sel}_{p}(E/L)^{G}=0$. The result then follows from Lemma 2.2, which asserts that $\operatorname{Sel}_{p}(E/L)=0$. ∎
@@
  • $v$ is unramified in $L$ and $E$ has nonsplit multiplicative,
  • $v$ splits completely in $L$ and $E$ has additive reduction at $v$,
@@
It follows from the definition of $S$ that $v$ must be totally ramified in $L$. Recall that Assumption 2.1 requires that $\widetilde{E}(k_{v})[p]=0$. The result then follows from [pathak2024rank, Proposition 2.8]. ∎
@@
First we consider the case in which $E$ has nonsplit multiplicative reduction at $v$. In this case, the result follows from [brau2014selmer, Proposition 5.8]. On the other hand, consider the case in which $E$ has split multiplicative reduction. Note that in light of Assumption 2.1, $v$ must be unramified in $L$ and $p\nmid c_{v}(E/K)$. Then, the result follows from [brau2014selmer, Proposition 5.7, (ii)]. Finally, consider the case in which $E$ has additive reduction at $v$. In this case, it is assumed that $v$ splits completely in $L$. In this case, $L_{w}=K_{v}$ for any prime $w$ which lies above $v$, and it is clear that $\gamma_{v}$ is injective. ∎
@@
  • First, assume that $v$ is unramified in $L$ and that $E$ has good ordinary reduction at $v$. In this case, $\gamma_{v}$ is injective by [brau2014selmer, Proposition 5.9, (ii)].
  • Next consider the case when $v$ is ramified in $L$ and $E$ has good ordinary reduction at $v$. In this case, it has been assumed that $\widetilde{E}(k_{v})[p]=0$, by Assumption 2.1. In this case, $\gamma_{v}$ is injective, according to [brau2014selmer, Proposition 5.9, (i)].
  • Consider the case when $v$ is unramified in $L$ and $E$ has good supersingular reduction at $v$. Then, $\gamma_{v}$ is injective by [brau2014selmer, Proposition 5.10, (ii)].
@@ 3. Asymptotics for elliptic curves
$h_{A,B}=h\left(E_{A,B}\right)\mathrel{\mathop{\ordinarycolon}}=\operatorname{max}\{|A|^{3},|B|^{2}\}.$
@@
$\mathcal{C}(X)\mathrel{\mathop{\ordinarycolon}}=\{(A,B)\in\mathcal{C}\mathrel{\mathop{\ordinarycolon}}h\left(E_{A,B}\right)\leq X\}.$
@@
Let $\mathcal{C}(X)$ be given as above. One has the following asymptotic estimate for the growth of $\mathcal{C}(X)$, as $X$ goes to $\infty$:
$\mathcal{C}(X)=\frac{4X^{5/6}}{\zeta(10)}+O\left(\sqrt{X}\right).$
@@
Any set of isomorphism classes of elliptic curves over $\mathbb{Q}$ can be identified with a subset $S$ of $\mathcal{C}$. Given $X>0$, set $S(X)\mathrel{\mathop{\ordinarycolon}}=S\cap\mathcal{C}(X)$. The density of $S$ (if it exists) is defined as follows
(3.1) $\mathfrak{d}(S)\mathrel{\mathop{\ordinarycolon}}=\lim_{X\rightarrow\infty}\frac{\#S(X)}{\#\mathcal{C}(X)}.$
When the limit in the above expression is replaced by $\limsup$ (resp $\liminf$), we write $\overline{\mathfrak{d}}(S)$ (resp. $\underline{\mathfrak{d}}(S)$) to denote the upper (resp. lower) density. The set $S$ is said to have positive density if $\underline{\mathfrak{d}}(S)>0$.
@@
The density of elliptic curves $E_{/\mathbb{Q}}$ with rank 0 (resp. $1$) is $1/2$ (resp. $1/2$).
The conjecture thus also predicts that the density of elliptic curves with rank $\geq 2$ is 0. Significant progress was made in seminal work of Bhargava and Shankar [BhargavaShankarbinary, bhargavashankar15, bhargavashankar5]. The idea is to relate Selmer groups of elliptic curves to integral forms. The integral forms are lattice points in certain fundamental regions in Euclidean space and techniques from the geometry of numbers are applied to study their distribution. Given a prime $p$, it is expected that the average size of the $p$-Selmer group of $E_{/\mathbb{Q}}$ is $(p+1)$. This has been proven to be the case for primes $p\leq 5$. We refer to work of Poonen and Rains [poonenrains] for predictions on the average size of Selmer groups and Tate–Shafarevich groups. More recently, Smith has proven various distribution results for the $2$-primary Selmer groups of twist families of abelian varieties. For further details, we refer to [Smith1, Smith2, smith3].
In this section, we fix a Galois extension $L/\mathbb{Q}$ with Galois group $\operatorname{Gal}(L/\mathbb{Q})\simeq\mathbb{Z}/p\mathbb{Z}$, where $p\in\{3,5\}$. Let $E_{/\mathbb{Q}}$ be an elliptic curve. Following Mazur and Rubin [Mazurrubindiophantine], $E$ is said to be diophantine stable in $L$ if the natural inclusion yields an equality $E(\mathbb{Q})=E(L)$. We assume throughout that $2$ and $p$ are unramified in $L$, and show that there is an effective positive density of elliptic curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$. This is done by showing that